# Modeling Language Evolution Using a Spin Glass Approach

### v1

@@ Abstract
Most traditional approaches to study diachronic language dynamics emphasize lexical data, but it would seem that a crucial contribution to the effectiveness of a thought-coding device is given by its core structure, its syntax. Based on the reduction of syntax to a set of binary syntactic parameters, we introduce here a model of natural language change in which diachronic dynamics are mediated by disordered interactions between/among parameters, even in the idealized limit of identical external inputs. We show in which region of ‘phase space’ such dynamics show the glassy features that are observed in natural language across time. In particular, binary syntactic vectors remain trapped in glassy metastable (tendentially stable) states when the degree of asymmetry in the disordered interactions is below a critical value, consistent with studies of spin glasses with asymmetric interactions. We further show that an added Hopfield-type memory term, would indeed, if strong enough, stabilize syntactic configurations even above the critical value, but losing the multiplicity of stable states. Finally, using a notion of linguistic distance in syntactic state space we show that a phylogenetic signal may remain among related languages, despite their gradually divergent syntax, exactly as recently pointed out for real-world languages. These statistical results appear to generalize beyond the dataset of 94 syntactic parameters across 58 languages, used in this study.
@@ I Introduction
The study of languages is crucial for a profound comprehension of human cognition, culture, and history. Yet, the most basic observation, that natural languages are diverse, remains theoretically unexplained. As the myth of the Tower of Babel indicates, it would make sense for an effective means of communication to be common to all those who might need to share their mental states. The counter-argument that some populations would benefit from their language being unintelligible to others seems overall weak. Further, it would seem plausible that structurally different language forms would significantly differ in their efficiency, leading human groups to discard the less efficient ones. Yet, structurally very different natural languages survive and thrive to this day. Rather than converging under selective pressures towards a common standard, they even appear to diverge, albeit very slowly, as if drifting away from each other. Drift dynamics might be expected to characterize superficial language features, such as the lexicon, but less so the internal syntactic structures – which must indeed be more resistant to change [1] but still in the direction of increased diversity.
@@
For two centuries, the study of language evolution has in fact largely relied on the form of lexical items [2]. Recently, Bayesian phylogenetic methods have been applied to core vocabulary [3], and geographic data has been integrated to model, e.g., the spread of Indo-European languages [4]. Historical constraints have been used to evaluate hypotheses on their place of origin [5], and automatic word comparison methods have enhanced reproducibility and bridged traditional approaches with machine learning [6].
@@
Studying the lexicon, however, can contribute little to address the basic questions above, because words can change in a myriad ways, and it is not feasible to quantify and contrast any potential change in language efficiency with the number of possibilities effectively sampled, what in a physical system would be the energy-entropy balance. Syntax, the structuring of words into sentences, is much more constrained: parametric approaches claim to reduce its diversity to a simple string of binary parameter values associated with each natural language [7]. The number of such parameters, initially thought to be few [8], has recently grown into the hundreds [9], but they still span a much more limited set than the one spanned by lexical items [10].
@@
In recent decades, in fact, phylogenetic linguistics has advanced through the application of quantitative methods to generative syntax and the comparison of parameter values [11, 12, 13, 14, 15, 16, 17, 9]. It has been shown that, contrary to long-held assumptions, syntactic features —- when analyzed within the generative biolinguistic framework —- can offer valuable insights into the long-term evolution of languages [17]. This approach reveals that syntax can carry a historical signal comparable to that of traditional etymological methods, but with sufficient verticality and greater chronological depth [17]. Recent research [9] has proposed that the integration of the Parametric Comparison Method (PCM) [14, 15, 17] with theories of possible languages [10] can elucidate phylogenetic relationships between language families. Following this line one may thus be able to extend the temporal scope of traditional comparative methods, offering a promising solution to long standing challenges in taxonomic linguistics [9].
@@
Chomsky’s original Principles &\&& Parameters model describes grammatical diversity by attributing cross-linguistic variation to binary syntactic parameters. It posits a universal and innate syntactic structure, with differences among languages arising from distinct parameter settings [18, 7], which are rapidly acquired during language learning by each individual. Language acquisition, in this view, thus involves selecting from a finite set of options [19]. More recent approaches within the minimalist program question the innateness of parameters, suggesting instead that they emerge dynamically through cognitive processes [20]. Still, aside from the innateness debate, linguistic variation is seen as governed by structured binary choices, offering a systematic account of phenomena such as variability in subject or object omission or word order.
@@
Syntactic parameter values cannot all be set, each independently of the others; parameters are interconnected through a hierarchical network of asymmetric implicational dependencies, exhibiting a near-logical structure shaped by functional constraints. For example, parameter B may be defined only if parameter A takes value +1. Seen the other way around, through the implications the value of one parameter may condition others. These can be regarded as strong, unidirectional interactions among parameters, that proceed always downstream from the top parameters, those that can be set independently of any other, toward the bottom of the hierarchy, occupied by parameters that can vary only subject to many conditions. In a network model, such dependencies may be formalized through polysynaptic weights, as detailed below.
@@
Unlike biological evolution, where adaptive pressures can usually be identified, that drive living systems toward greater efficiency (or fitness), syntactic change does not appear to follow such a straightforward route. While historical syntactic states can often be reconstructed, the evolution of syntax - lacking clear functional directions - remains both constrained and unpredictable. Its trajectories are marked by two seemingly contradictory features: persistence, where certain structures remain stable over time, and diversification, where languages diverge into distinct syntactic systems.
@@
The number and the complex interactions among syntactic parameters make statistical methods a reasonable option for their study. Although parameter values likely vary in subtle ways also across individuals nominally speaking the same natural language, the sets of parameter values of natural languages can be regarded as so-called metastable states. That is, they are attractive – attractive to all the individuals sharing a particular language – and not absolutely but largely stable, over centuries. The diversity expressed by the observed sets can be interpreted as stemming from the lack of any overarching order or symmetry principle, which might have imposed common values. This situation parallels glassy dynamics, in which numerous weak, disorderly interactions among components produce persistent, quasi-stable states that do not appear to conform to any ordering or symmetry principle. Therefore, disorder may play a pivotal role in syntactic evolution by enabling both diversification and persistence within language systems.
@@
It should be emphasized that these properties appear rather trivial if considered individually: diversification – without attractor dynamics – can be explained away by different external inputs, while persistence – if diachronic changes were negligible – is obviously a generic property of static systems. To analyze, however, how these properties can coexist and interact, it seems necessary to go beyond linear reasoning and adopt a quantitative model.
@@
In the following, in Section II, we introduce a novel model of syntactic change inspired by glassy systems. This model combines strong, explicitly modeled asymmetric interactions that tendentially produce implicational relations, with a multitude of weak, statistically modeled pairwise interactions, expressing disorder in the exact way parameters influence each other. The model can be used to investigate the overall impact of these effectively random couplings on system dynamics. Section III presents a sharp transition obtained by slightly varying an angle defining the average asymmetry in the disordered interactions, and locates the transition point. In Section IV, we incorporate a Hebbian term as an additional symmetric interaction weight between parameters, and examine its effect on evolutionary dynamics. Finally, in Section V, we focus on language distances, analyzing the ‘historical signal’, which reflects linguistic relatedness.
@@
Figure 1: Symmetric random couplings can block syntax evolution. Simulations illustrating parameter fluidity as a function of implication strength $\zeta$ and of the degree of symmetry $\phi$ in the random interactions. The black datapoints are the number of flipping parameters as a function of time for $\zeta\to\infty$, i.e., with implications alone. Syntactic structures then continue to evolve, with 20–30 parameters flipping at each time step, indicating fluid dynamics. When adding random interactions, the strength $\zeta=0.01,0.1,1,10,100,1000$ of the implications is denoted by progressively darker colors. With random symmetric connections, ($\phi=0^{\circ}$, panel a), after the first few time steps, parameters stop changing, whatever the value of $\zeta$. Eventually, after more time steps, the same freezing occurs in panel (b), with $\phi=30^{\circ}$; with more asymmetry, instead, i.e., $\phi=60^{\circ}$, (c) and $\phi=90^{\circ}$, (d), fluidity is maintained throughout, with the parameters continuing to flip over time, actually the more the weaker are the implications. Each of 58 natural languages is used as a starting point for 10 simulations, for the same set of weak interaction weights; and then averages are taken over 100 distinct sets of weights. Therefore, the curves shown are averages over a total of 58,000 data points.
@@ II Modeling Syntactic Evolution
This study is based on simulating a simple model of the evolution of syntax over time. The model assumes that parameters can change value (or flip) under the influence of other parameters, and with negligible external inputs. Language change therefore arises from the attempt, by the parameters, to coordinate their values, in what is effectively a recurrent network. The crucial assumption is that complete coordination may never be possible, given the lack of intentional design: a large array of disorderly constraints tends to leave those mutually incompatible, to some extent, and prone to change. The assumption of negligible external inputs does not imply that we believe them to be irrelevant to real syntax dynamics, of course; it is simply a methodological choice adopted with models of other complex systems addressing the same critical issue [21]: do they allow for a multiplicity of attractor states even when isolated from external inputs?
@@
Internal influences are of two distinct types: implications, which explicitly reproduce linguistically derived conditions on parameter values, and are inherently asymmetric; and random couplings, which represent all presumable interactions among parameters that have not yet been analyzed explicitly, and which therefore we treat statistically. They are randomly assigned in this statistical sense, but we assume them to be the same across all languages, and keep them fixed through their evolution. Implications may involve multiple conditions, e.g. parameter $i$ can be defined and hence may take its non-default value only if either parameter $j$ or parameter $k$ is set to +1. Their overall strength (relative to the random couplings) is scaled by a coefficient $\zeta$, which implies that for $\zeta\to\infty$ implications become absolute rules, while for $\zeta\to 0$ they reduce to weak biases. The random couplings, instead, reflect the assumption that all parameters, implemented within one brain, must perforce interact in a multitude of ways not individually analyzable: to hope that they could be isolated from each other appears to be wishful thinking. In this study, we consider only the simplest interactions, i.e. pairwise ones, postponing more complex variants to future analyses. We introduce an angle $\phi$ that statistically quantifies the degree of symmetry of such random couplings, from $\phi=0^{\circ}$ (fully symmetric couplings, e.g., from parameter $j$ to parameter $i$ and viceversa the two coupling coefficients would be equal, $W_{ij}=W_{ji}$); to $\phi=90^{\circ}$ (anti-symmetric couplings, i.e., they would be opposite, $W_{ij}=-W_{ji}$).
@@
The model is applied to the most extensive available corpus of parametric syntax: the PCM dataset put together by Longobardi and collaborators [14, 9]. It includes $L=58$ languages from various families, each characterized by $N=94$ syntactic parameters relative to the domain of nominal syntax. Each of the $58\times 94$ parameter values, denoted as $\{S_{i}^{\mu}\}$, with $i=1,\ldots,N$ and $\mu=1,\ldots,L$, is either set to $\{S_{i}^{\mu}\}=-1$ (default) or to $\{S_{i}^{\mu}\}=+1$ (the marked value). Note that our notation aims to bridge with the physics literature on spin systems, hence binary values are identified with Spins [22]; and with that on associative memories, hence distinct language configurations are indexed by the symbol $\mu$ [21]. We refrain from using the symbol $p$ for either the number of parameters considered or, as in the associative memory literature, the number of distinct patterns or configurations of values they are supposed to be able to express, and use $N$ and $L$ instead, respectively.
@@
In the PCM database, many parameters can also take the value 0, or undefined, if they are made undefined or irrelevant by the values of other parameters hierarchically above them. Here, however, we identify the undefined with the default state. This is a simplification which may not be generally valid, and which we plan to remove in future studies. With such a simplification, each language is then assigned a binary list of syntactic parameter values. The implications already limit the number of possible combinations the latter can take. Our goal is to study how the additional random couplings further affect changes in time of the binary vector associated with a community of speakers.

### v2

@@ Abstract
Most traditional approaches to study diachronic language dynamics emphasize lexical data, but it would seem that a crucial contribution to the effectiveness of a thought-coding device is given by its core generative structure, i.e., its syntax. Based on the reduction of syntax to a set of binary syntactic parameters, we introduce here a model of natural language change in which diachronic dynamics stem from disordered interactions between/among parameters, even in the idealized limit of identical external inputs. We show in which region of “phase space” such dynamics show the glassy features that are observed in natural language across time. In particular, binary syntactic vectors remain trapped in glassy metastable (i.e., tendentially stable) states when the degree of asymmetry in the disordered interactions is below a critical value, consistent with studies of spin glasses with asymmetric interactions. We further show that an added Hopfield-type memory term, would indeed, if strong enough, stabilize syntactic configurations even above the critical value, but losing the multiplicity of stable states. Finally, using a notion of linguistic distance in syntactic state space we show that a phylogenetic signal may remain among related languages, despite their gradually divergent syntax, exactly as recently pointed out for real-world languages. These statistical results appear to generalize beyond the dataset of 94 syntactic parameters across 58 languages used in this study.
@@ I Introduction
The study of languages is crucial for a profound comprehension of human cognition, culture, and history. Yet, the most basic observation, that natural languages are diverse, remains theoretically unexplained. As the myth of the Tower of Babel indicates, it would make sense for an effective means of coding thoughts to be common to all those who might need to share their mental states. Further, it would seem plausible that structurally different language forms would significantly differ in their efficiency, leading human groups to discard the less efficient ones. Yet, structurally very different natural languages survive and thrive to this day. Rather than converging under selective pressures towards a common standard, they even appear to diverge, albeit very slowly, as if drifting away from each other [guardiano2005parametric, ceolin2021boundaries]. Drift dynamics might be expected to characterize superficial language features, such as the lexicon, but less so the internal syntactic structures. These structures are indeed more resistant to change [keenan2014universal], but they still change in the direction of increased diversity.
@@
For two centuries, the study of language evolution has in fact largely relied on the form of lexical items [swadesh1952lexico]. Recently, Bayesian phylogenetic methods have been applied to core vocabulary [gray2003language, sagart2019dated, greenhill2020bayesian], and geographic data has been integrated to model, e.g., the spread of Indo-European languages [bouckaert2012mapping]. Historical constraints have been used to evaluate hypotheses on their place of origin [chang2015ancestry], and automatic word comparison methods have enhanced reproducibility and bridged traditional approaches with machine learning [list2017potential].
@@
Studying the lexicon, however, can contribute little to address the basic questions above, because words can change in a myriad ways, and it is not feasible to quantify and contrast any potential change in language efficiency with the number of possibilities effectively sampled, what in a physical system would be the energy-entropy balance. Furthermore, the study of lexical evolution has well known space-time limits defined by the concept of language family. Syntax, the structuring of words into sentences, is much more constrained: parametric approaches claim to reduce its diversity to a simple string of binary parameter values associated with each natural language [chomsky1993theory]. The number of such parameters, initially thought to be few [baker2008atoms], has recently grown into the hundreds [ceolin2021boundaries], but they still span a much more limited set than the one spanned by lexical items [moro2016impossible].
@@
In recent decades, in fact, phylogenetic linguistics has advanced through the application of quantitative methods to generative syntax and the comparison of parameter values [roberts2007diachronic, biberauer2005changing, longobardi2008syntactic, longobardi2009evidence, longobardi2013toward, galves2012parameter, ceolin2020formal, ceolin2021boundaries]. It has been shown that, contrary to long-held assumptions, syntactic features —- when analyzed within the generative biolinguistic framework —- can offer valuable insights into the long-term evolution of languages [ceolin2020formal]. This approach reveals that syntax can carry a historical signal comparable to that of traditional etymological methods, but with sufficient verticality and greater chronological depth [ceolin2020formal]. Recent research [ceolin2021boundaries] has proposed that the integration of the Parametric Comparison Method (PCM) [longobardi2009evidence, longobardi2013toward, ceolin2020formal] with theories of possible languages [moro2016impossible] can elucidate phylogenetic relationships between language families. Following this line one may thus be able to extend the temporal scope of traditional comparative methods, offering a promising solution to long standing challenges in taxonomic linguistics [ceolin2021boundaries].
@@
Chomsky’s original Principles &\& Parameters model describes grammatical diversity by attributing cross-linguistic variation to binary syntactic parameters. It posits a universal and innate syntactic structure, with differences among languages arising from distinct parameter settings [chomsky1993lectures, chomsky1993theory], which are rapidly acquired during language learning by each individual. Language acquisition, in this view, thus involves selecting from a finite set of options [fodor1998unambiguous]. More recent approaches within the minimalist program question the innateness of parameters, suggesting instead that they emerge dynamically through cognitive processes [karimi2017introduction]. Still, aside from the innateness debate, linguistic variation is seen as governed by structured binary choices, offering a systematic account of phenomena such as variability in subject or object omission or word order.
@@
Syntactic parameter values cannot all be set, each independently of the others; parameters are interconnected through a hierarchical network of asymmetric implicational dependencies, exhibiting a near-logical structure shaped by functional constraints [baker2008atoms, guardiano2016parameter]. For example, parameter B may be defined only if parameter A takes value +1. Seen the other way around, through the implications the value of one parameter may condition others. These can be regarded as strong, unidirectional interactions among parameters, that proceed always downstream from the top parameters, those that can be set independently of any other, toward the bottom of the hierarchy, occupied by parameters that can vary only subject to many conditions (see Fig.A.1, which refers to the parameters and their full implicational structure shown in Table S2 of [ceolin2021boundaries]). In a network model, such dependencies may be formalized through polysynaptic weights, as detailed below.
@@
Unlike biological evolution, where adaptive pressures can usually be identified that drive living systems toward greater efficiency (or fitness), syntactic change does not appear to follow such a straightforward route. While past syntactic states can often be reconstructed, the evolution of syntax – lacking clear functional directions – remains constrained but unpredictable. Its trajectories are marked by two seemingly contradictory features: persistence, where certain structures remain stable over time, and diversification, where languages diverge into distinct syntactic systems.
@@
The number and the complex interactions among syntactic parameters make statistical methods a reasonable option for their study. Although parameter values likely vary in subtle ways also across individuals nominally speaking the same natural language, the sets of parameter values of natural languages can be regarded as so-called metastable states. That is, they are attractive – attractive to all the individuals sharing a particular language – and not absolutely but largely stable, over centuries. The diversity expressed by the observed sets can be interpreted as stemming from the lack of any overarching order or functional principle, which might have imposed common values. This situation parallels glassy dynamics, in which numerous weak, disorderly interactions among components produce persistent, quasi-stable states that do not appear to conform to any ordering or symmetry principle. Therefore, disorder may play a pivotal role in syntactic evolution by enabling both diversification and persistence within language systems.
@@
It should be emphasized that these properties appear rather trivial if considered individually: diversification – without attractor dynamics – can be explained away by different external inputs, while persistence – if diachronic changes were negligible – is obviously a generic property of static systems. If syntax were, however, so directly sensitive to external inputs as required to produce diversification on its own, it would remain to be seen where it could find the stability necessary for speakers to share a grammar. In the most plausible case the relation between some hard-to-define primitive source in the external input and actual syntactic change is mediated by a number of internal factors: hence the necessity for first studying the overall functioning of such factors. This is an issue that has been avoided by theories of languages as synchronic codes or sets of rules, but has to be confronted by perspectives that view them as dynamical systems. To analyze, therefore, how diversification and persistence can coexist and interact, it seems necessary to go beyond linear reasoning and adopt a quantitative, multifactorial dynamical model.
@@
In the following, in Section II, we introduce a novel model of syntactic change inspired by glassy systems. This model combines strong, explicitly modeled asymmetric interactions that tendentially produce implicational relations, with a multitude of weak, statistically modeled pairwise interactions, expressing disorder in the exact way parameters influence each other. Already in its vanilla version, with purely binary parameters, the model can be used to probe the overall impact of these effectively random couplings on system dynamics. Section III presents a sharp transition obtained by slightly varying an angle defining the average asymmetry in the disordered interactions, and locates the transition point. In Section IV, we incorporate a Hebbian term as an additional symmetric interaction weight between parameters, and examine its effect on evolutionary dynamics. In Section V, we focus on language distances, analyzing the ‘historical signal’, which reflects linguistic relatedness, Finally, in Section VI, we show that the above findings extend to a refined version of the model, in which implications determine whether a parameter can be defined or not, given the values of those above it in the hierarchy. Conclusions and future perspectives are summarized in Section VII.
@@
Figure 1: Symmetric hidden couplings can block syntax evolution. Simulations illustrating parameter fluidity as a function of implication strength $\zeta$ and of the degree of symmetry $\phi$ in the hidden interactions. The black datapoints are the number of flipping parameters as a function of time for $\zeta\to\infty$, i.e., with implications alone. Syntactic structures then continue to evolve, with 20–30 parameters flipping at each time step, indicating fluid dynamics. When adding hidden interactions, the strength $\zeta=0.01,0.1,1,10,100,1000$ of the implications is denoted by progressively darker colors. With hidden symmetric connections, ($\phi=0^{\circ}$, panel a), after the first few time steps, parameters stop changing, whatever the value of $\zeta$. Eventually, after more time steps, the same freezing occurs in panel (b), with $\phi=30^{\circ}$; with more asymmetry, instead, i.e., $\phi=60^{\circ}$, (c) and $\phi=90^{\circ}$, (d), fluidity is maintained throughout, with the parameters continuing to flip over time, actually the more the weaker are the implications. What stabilizes very quickly is the average number of flipping parameter per time step. Each of 58 natural languages is used as a starting point for 10 simulations, for the same set of weak interaction weights; and then averages are taken over 100 distinct sets of weights. Therefore, the curves shown are averages over a total of 58,000 data points.
@@ II Modeling Syntactic Evolution
This study is based on simulating a simple model of the evolution of syntax over time. The model assumes that parameters can change value (or flip) under the influence of other parameters, and with negligible external inputs. Language change therefore proceeds through the attempt, by the parameters, to coordinate their values, in what is effectively a recurrent network. The crucial assumption is that complete coordination may never be possible, given the lack of intentional design: a large array of disorderly constraints tends to leave those mutually incompatible, to some extent, and prone to change. The assumption of negligible external inputs does not imply that we believe them to be irrelevant to real syntax dynamics, of course a multifactorial process; it is simply a methodological choice adopted with models of other complex systems addressing the same critical issue [amit1989modeling]: do they allow for a multiplicity of attractor states even when isolated from external inputs?
@@
Internal influences may be of two distinct types: the implications cited above and hidden couplings, which represent all presumable interactions among parameters that have not yet been analyzed explicitly, and which therefore we treat statistically. The assumption is that parameters, as they are somehow implemented within the same brain, cannot be completely isolated from each other and so must interact in a multitude of ways, which we do not know. The hidden couplings are randomly assigned in this statistical sense, but we assume them to be the same across all languages, as they should reflect common neural mechanisms, and we keep them fixed as syntax evolves in our simulations. We further assume in this study that each hidden coupling $W_{ij}$ involves only a pair of parameters, the influencer, $j$, and the influenced, $i$; a separate coupling $W_{ik}$ expresses the influence of another parameter $k$ on $i$, and yet another one $W_{ji}$ of $i$ on $j$. Non-pairwise or more complex hidden interactions are left for future studies. Importantly, we introduce an angle $\phi$ that statistically quantifies the degree of symmetry of such hidden couplings, from $\phi=0^{\circ}$ (fully symmetric couplings, e.g., from parameter $j$ to parameter $i$ and viceversa the two coupling coefficients would be equal, $W_{ij}=W_{ji}$); to $\phi=90^{\circ}$ (anti-symmetric couplings, i.e., they would be opposite, $W_{ij}=-W_{ji}$).
Implications, on the other hand, explicitly reproduce linguistically derived conditions on parameter values, are inherently asymmetric and may involve multiple influencing parameters in a complex combination, e.g. parameter $i$ can be defined and hence may take its non-default value only if either parameter $j$ or parameter $k$ is set to +1.
@@
The model is applied to the most extensive available corpus of parametric syntax: the PCM dataset put together by Longobardi and collaborators [longobardi2009evidence, ceolin2021boundaries]. It includes $L=58$ languages from various families, each characterized by $N=94$ syntactic parameters relative to the domain of nominal syntax. Each parameter value, denoted as $\{S_{i}^{\mu}\}$, with $i=1,\ldots,N$ and $\mu=1,\ldots,L$, is set, if the parameter can be defined in that language, to either $\{S_{i}^{\mu}\}=-1$ (default) or to $\{S_{i}^{\mu}\}=+1$ (the marked value). Note that our notation aims to bridge with the physics literature on spin systems, hence binary values are identified with Spins [sherrington1975solvable]; and with that on associative memories, hence distinct language configurations are indexed by the symbol $\mu$ [amit1989modeling]. We refrain from using the symbol $p$ for either the number of parameters considered or, as in the associative memory literature, the number of distinct patterns or configurations of values (==languages) they are supposed to be able to express, and use $N$ and $L$ instead, respectively.
@@
In the PCM database, many parameters can also take the value 0, or undefined, if they are made undefined or irrelevant by the values of other parameters hierarchically above them. Here, we approach the issue of the undefined state in two different ways, as detailed below and in Sect.VI. Each language is then assigned a list of syntactic parameter values. The implications already limit the number of possible combinations the latter can take. Our goal is to study how the additional hidden couplings further affect changes in time of the vector of values associated with a community of speakers.
II.1 The purely binary model
In a first, basic version of our model, parameters can take only the marked value +1+1 or not take it, then remaining in the default state 1-1. In a later version, considered in Sect.VI, the unmarked value 1-1 is distinct from the undefined condition 0, and parameters exert influences through the hidden couplings only when they are defined, in states +1+1 or 1-1.
In this first version, implications and interactions mediated by hidden couplings add up in the same equations, and the overall strength of the implications (relative to the hidden couplings) is scaled by a coefficient $\zeta$, which implies that for $\zeta\to\infty$ implications become absolute rules, while for $\zeta\to 0$ they reduce to weak biases.